aenglebright:

An athlete is running in a marathon that eventually finishes at a local high school. After x minutes, the distance the athlete is from the school is given by y (in miles), where x + 10y = 260 (Note: 0 ≤ x ≤ 10 and 0 ≤ y ≤ 26). Complete the steps below to understand how to find the intercepts of the equation and what they represent.

TheSmartOne:

Hey there! Welcome to QuestionCove!! Can you post the complete question? Also something doesn't seem right if x is bound between 0 and 10, because if the total marathon is 26 miles, idek But please post the rest of the question :)

aenglebright:

An athlete is running in a marathon that eventually finishes at a local high school. After x minutes, the distance the athlete is from the school is given by y (in miles), where x + 10y = 260 (Note: 0 ≤ x ≤ 10 and 0 ≤ y ≤ 26). Complete the steps below to understand how to find the intercepts of the equation and what they represent. Use the relationship between x and y to find the time it takes for the runner to reach the school. Use the relationship between x and y to find the distance between the runner and the high school at the start line. If the runner ran for 40 minutes from the start line, what is the remaining distance between the runner and the school?

TheSmartOne:

Perfect, so do you know how to find the x-intercept or y-intercept?

aenglebright:

x is minutes and y is miles

TheSmartOne:

Hint: To find the x-intercept, set y=0 and solve for x To find the y-intercept, set x=0 and solve for y

TheSmartOne:

Yes, what you listed is the variables and what they represent in our equation the x-intercept is when the line crosses the x-axis and this happens when y=0 And vice versa for the y-intercept, it's when it crosses the y-axis and that happens when x=0

TheSmartOne:

So x + 10y = 260 To find the x-intercept Plug in y=0 and solve for x To find the y-intercept Plug in x=0 and solve for y

aenglebright:

ok

aenglebright:

260 represents?

TheSmartOne:

so think about it, these intercepts are like the end points for our line earlier you said that y was miles, so when we're 0 miles away from the school that means we've run the whole marathon, and you get x = 260 so that means... How long did it take you to run the marathon?

TheSmartOne:

(they're the end points because we can't go into negatives or anything, like we can't keep running after we reach the school, or it's not possible to be running away from the school) just to clarify

aenglebright:

26 miles

TheSmartOne:

Yes so basically at x=0, we get y=260 So when we have run 0 minutes, we're at the beginning of the marathon, we're 26 miles away And x=260, y=0 We have run for 260 minutes, and then we'll finally be 0 miles away from the school so it'll take 260 minutes to do the 26 miles

aenglebright:

If the runner ran for 40 minutes from the start line, what is the remaining distance between the runner and the school?

TheSmartOne:

So he ran for 40 minutes, Just plug in x=40 into your equation and solve for y Also this is why the 0 ≤ x ≤ 10 is a typo because it's not possible to limit ourselves to 10 minutes

TheSmartOne:

but ye x + 10y = 260 plug in x=40 and then the number you get for y is how many miles are left to run

aenglebright:

22

TheSmartOne:

Bingo!!

aenglebright:

How much time will the runner take to complete half the distance between the start line and the high school?

TheSmartOne:

So if the total distance is 26 miles, how much would half of that be? Plug that in to y and solve for x

aenglebright:

x+13=260

TheSmartOne:

It's 10y, so you'll have to multiply by 10. Don't forget that

aenglebright:

is the answer 20

TheSmartOne:

not quite, how did you get to that? x + 10y = 260 plug in y=13

aenglebright:

x= 130 minutes.

TheSmartOne:

Good job!

aenglebright:

can you help me with another question? Jason works as an administrative assistant for a local lawyer's office. The office is running low on pens and notepads, so he is asked to buy $150 worth of pens and notepads. Pens cost $1, and notepads cost $2. If the number of pens he buys is represented by x and the number of notepads he buys is represented by y, the situation can be given by x + 2y = 150, where 0 ≤ x ≤ 150, 0 ≤ y ≤ 75 and x and y can take only integer values. Assuming that the stationery store is out of notepads, how many pens would Jason have to purchase to spend the entire amount he is given? Assuming that the stationery store is out of pens, how many notepads would Jason have to purchase to spend the entire amount he is given?